OpenStudy (anonymous):

the expression 3x-y-[2x+3(2-y)] is equal to what?

OpenStudy (anonymous):

3x-y-[2x+3(2-y)] =3x-y-(2x+6-3y) = 3x-y-2x-6+3y =3x-2x+3y-y = x+2y

OpenStudy (anonymous):

but the choices are x-4y-6, x-6, x+2y-6,x-4y+6

OpenStudy (anonymous):

o o i missed something the answer is x+2y-6

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

3x-y-[2x+3(2-y)] =3x-y-(2x+6-3y) = 3x-y-2x-6+3y = 3x-2x+3y-y-6 = x+2y-6

OpenStudy (anonymous):

oh i get it! :) thanks :)

OpenStudy (anonymous):

can you help me with other problems? hehe

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ x ^{3}-3x ^{2}+7 is subtracted from 5x ^{3}-4x ^{2}-7\] is equal to?

OpenStudy (anonymous):

( x^3 -3x^2 +7 ) - ( 5x^3 -4x^2 - 7 ) = x^3 -3x^2 +7 - 5x^3 +4x^2 +7 = x^3-5x^3 -3x^2 -4x^2 +7+7 do the last step now

OpenStudy (anonymous):

\[4^{3}-x^{2}-14 \] is this right?

OpenStudy (anonymous):

* 4x^3